In playing Monopoly, rolling doubles three times in a row sends you to jail. If you roll the dice 100 times over the course of a game, what is the probability that you will have rolled doubles, at any point, exactly three times
Answer by stanbon(75887)

In playing Monopoly, rolling doubles three times in a row sends you to jail. If you roll the dice 100 times over the course of a game, what is the probability that you will have rolled doubles, at any point, exactly three times
--------------------------------
Binomial Problem with n = 100 and p(doubles) = 1/6
-----
P(x = 3) = 100C3(1/6)^3*(5/6)^97 = binompdf(100,1/6,3) = 0.000156..
------------
Cheers,
Stan H.
===================